A rancher sells hides and beef. The two goods are assumed to be jointly produced in fixed
proportions. The marginal cost equation for the beef-hide product package is given by
MC = 30 + 5Q
The demand and marginal revenue equations for the two products are:
BEEF
HIDES
P = 60 – 1Q
P = 80 – 4Q
MR = 60 – 2Q
MR = 80 – 4Q
What prices should be charged for beef and hides? How many units of the product package
should be produced?
1
Expert's answer
2021-09-24T19:03:31-0400

Solution:

First, determine the output to be produced:

MR = MC

MR for Beef = 60 – 2Q

MR for Hides = 80 – 4Q

Find the total MR fir both Hides and Beef:

60 – 2Q + 80 – 4Q = 140 – 6Q

MC = 30 + 5Q

140 – 6Q = 30 + 5Q

140 – 30 = 5Q + 6Q

110 = 11Q

Q = 10


Substitute Q in the demand functions to determine the prices:

Beef demand: 60 – 1Q

= 60 – 1(10) = 60 – 10 = 50

The price that should be charged for beef = 50


Hides demand: 80 – 2Q

= 80 – 2(10) = 80 – 20 = 60

The price that should be charged for Hides = 60

 

The units of the product package that should be produced = 10 units